OpenStudy (anonymous):

solve the equatuion sqare root of x+4=7

OpenStudy (saifoo.khan):

\[\sqrt{x+ 4} = 7\]

OpenStudy (saifoo.khan):

this?

OpenStudy (anonymous):

yess

OpenStudy (saifoo.khan):

x = 49-4 x = 45

OpenStudy (anonymous):

thankss big help